# Filtering Discomforting Recommendations with Large Language Models

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Personalized algorithms, which analyze user preferences to deliver tailored content and thereby support human decision-making, are indispensable across web platforms (Ricci et al., 2021; Konstan and Terveen, 2021). While these algorithms are designed to enhance user experience, they can inadvertently expose users to discomforting recommendations (Park et al., 2024). For example, if a user searches for sensitive topics like health issues, the algorithm might suggest related health products, which could be perceived as a breach of privacy and cause unease (Ur et al., 2012). Similarly, when a user is experiencing emotional distress, such as after a breakup, algorithms lacking contextual awareness might recommend content that evokes painful memories, potentially worsening the user’s emotional state (Pinter et al., 2019). Such recommendations may not only fail to engage users but also lead to negative emotional consequences, such as anxiety, unease, or distress (Park et al., 2024). The perception of discomfort is highly subjective, meaning that content one user finds enjoyable may be discomforting to another (Park et al., 2024; Smith et al., 2022; Pierce et al., 2018; Singh, 2019). This subjectivity underscores the urgent need for a more nuanced approach to discomforting content identification that aligns more closely with individual user experiences (Harris and Cook, 2011; Reichmann et al., 2023; Shahi and Kana Tsoplefack, 2022).
@@
Figure 1. Problem formulation.
Figure 2. The workflow of DiscomfortFilter.
@@ 2. Related Work
Personalized algorithms primarily derive user preferences from behavioral data, leading to incomplete user modeling and discomforting recommendations (Ricci et al., 2021). In Appendix E.1, we provide a detailed review of studies illustrating this phenomenon across various scenarios, including privacy invasion (Ur et al., 2012), lack of contextual understanding (Pinter et al., 2019), popularity bias (Ferraro, 2019), and information bubbles (Rowland, 2011), along with their negative outcomes (Smith et al., 2022). Although previous research has identified scenarios and causes of discomfort, few studies have focused on designing systems to identify and filter these recommendations, which our work aims to address.
In Appendix E.2, we provide a detailed review of two potential solutions for filtering discomforting recommendations: interactive recommendation systems (Lyons et al., 2021; He et al., 2016; Ge et al., 2022) and content moderation systems (Cresci et al., 2022; Gurkan et al., 2024; Jhaver et al., 2023). Both allow users to modify recommendations to mitigate discomfort. However, interactive systems often struggle to scale in opaque environments due to their reliance on specific algorithm designs, while moderation systems focus on objectively harmful content, which may be less effective for addressing subjective discomfort. Thus, the subjectivity of discomfort perception and the opacity of algorithms present notable challenges.
In Appendix E.3, we provide a detailed review of studies on LLMs as personal assistants (Li et al., 2024), covering commercial applications (Mehdi, 2023a), frameworks for human-centered recommender systems (Shu et al., 2024), and designs tailored for special populations (Devries et al., 2024). The impressive capabilities of LLMs, especially in handling personal data and services, underscore their potential for effectively filtering discomforting recommendations. To our knowledge, this is the first exploration of using LLMs for this specific purpose.
@@ 4.2.1. Content Filter Module
A user can manage filtering rules directly through this module. These rules are described in natural language, specifying the discomforting recommendations the user wishes to avoid. Each filtering rule has an associated activation option. During the filtering phase, the module reviews each recommendation against all active filtering rules to determine if the content matches any discomforting criteria. Only recommendations that do not match any filtering rules will be shown to the user; otherwise, they will be filtered out (G4). The identification process uses a chain-of-thought (CoT) method, with the prompt detailed in Appendix C.1. Filtering records will be saved and forwarded to the Filtering Needs Discovery Module.
@@ 4.2.2. Preference Profile Construction Module
The preference profile is stored and sent to the Filtering Needs Discovery Module to improve the personalization of the conversational agent and provide users with clear explanations. The prompt used for this multi-agent pipeline can be found in Appendix C.2.
@@ 5.1.4. Results
Figure 7. Score distribution from the questionnaire: the integers in the colored bars represent the number of participants for each score, with the color-to-score relationship indicated in the bottom-right corner.
@@ 5.2.2. Process
# Messages # Add # Update
Content Filter Module - 2.9 2.1
Strategy 1 13.0 4.4 2.7
Strategy 2 22.3 3.2 7.3
@@ Acknowledgements.
Sincere thanks to all participants involved in this work.
@@ Appendix B The Detailed Process of the User Study
Figure 10. The workflow of the Candidate Rule Generation Module.
@@ E.1. Discomforting Recommendations
Table 4. Confusion matrix.
Predicted Positive Predicted Negative
Actual Positive 173 32
Actual Negative 67 208
@@ E.2.2. Content Moderation Systems
Content moderation is used to regulate harmful content, such as hate speech and abuse (Alghowinem, 2019; Liu et al., 2023; Franco et al., 2023; Jhaver et al., 2019; Ye et al., 2023; Park et al., 2021). The process has evolved from manual review (Gillespie, 2018; Seering et al., 2019) to automated methods using NLP technologies (Chandrasekharan et al., 2019; Horta Ribeiro et al., 2023; Long et al., 2017), which has led to reduced transparency in moderation and less feedback for users (Wright et al., 2021; Jhaver et al., 2019). To address these challenges, RECAST (Wright et al., 2021) provides mechanisms for explanation, revision, and user feedback in the moderation process. Recent research on using LLMs for content moderation (Ma et al., 2023; Markov et al., 2023; Ye et al., 2023; Mullick et al., 2023) highlights their potential in tasks like toxicity detection (Kumar et al., 2023) and rule violation identification (Kolla et al., 2024). At the same time, there is increasing support for decentralizing content moderation, shifting control from centralized platforms to individual users, thus better enabling them to manage content they wish to avoid on social media (Jhaver et al., 2022). This shift recognizes that a one-size-fits-all approach to content regulation fails to meet the diverse needs of the user base (Jiang et al., 2023; Jhaver et al., 2023).
@@ E.3. LLMs as Personal Assistants
Several commercial products, such as Microsoft’s Copilot, New Bing, Google’s Bard, and Gemini, as well as offerings from smartphone manufacturers, have integrated LLMs to boost productivity and enhance web and on-device experiences (Mehdi, 2023a; Achiam et al., 2023; Mehdi, 2023b). In the area of human-centered recommender systems, the RAH framework (Shu et al., 2024), an LLM-based multi-agent system, analyzes user preferences by observing ratings and reviews and incorporates a reflection mechanism to improve the accuracy. Recent research has explored the use of LLMs as personal assistants for various special populations. Glasser et al. (2022) examined how deaf and hard-of-hearing users interact with LLM-based personal assistants capable of understanding sign language to better customize these technologies for their needs. Additionally, some work focuses on designing interfaces for visually impaired individuals to access visual information using LLMs and visual language models (Zhang, 2024). Finally, Jang et al. (2024) investigated how autistic workers use LLMs for communication assistance, highlighting their reliance on and trust in these tools.
@@ F.0.1. Generalizability to non-text-based items
In Section 6.2, we suggested future extensions to multimodal content. Expanding to non-text scenarios is a natural progression, requiring the use of multimodal models to understand such content. However, the efficiency of processing, especially for videos, remains a concern. Preprocessing non-text content offline may offer a viable solution.
@@ F.0.2. Differences between discomforting and disliked items
F.0.3. Discussion of critiquing-based recommender systems.
@@
F.0.4. Computational overhead and scalability.
@@
F.0.5. Discussion on the echo chamber problem.
Research indicates that merely providing users with diverse recommendations is not sufficient to avoid the echo chamber problem (Zhang et al., 2024). Active exploration by users is key to addressing the echo chamber effects (Zhang et al., 2024). We believe that, through actively guiding users, the critiquing-based recommender systems, as well as our work, can contribute to achieving this goal.
@@
F.0.6. It would be helpful to compare the effectiveness of the filtering approach presented in this paper with those used in other studies on recommendation unlearning.
Recommendation unlearning emphasizes the model’s capability, primarily focusing on the completeness, efficiency, and performance of the model after forgetting. These methods typically predefine which interactions should be forgotten, assess the completeness of forgetting through Membership Inference Attacks (MIA), measure the efficiency of forgetting through runtime analysis, and evaluate post-forgetting performance through recommendation accuracy. In contrast, DiscomfortFilter filters interactions from the perspective of user needs—due to its subjectivity, it is difficult to clearly define which interactions should be filtered in advance. Therefore, it is hard to make a quantitative comparison with recommendation unlearning, as current unlearning methods cannot handle subjective user inputs. We believe that these two types of work focus on model design and human-centered design, respectively, and are complementary to each other.
F.0.7. It is unclear how well this effect will work for ephemeral filtering criteria.
@@
F.0.8. Only the MIND dataset was used in the offline experiment.
As shown in Figure 5, DiscomfortFilter constructs positive and negative sample pairs by observing the content recommended by the algorithm and the user’s click behavior on that content. Specifically, our negative sampling is not performed randomly from the entire item set; instead, it is selected from items that appeared alongside the clicked content but were not clicked. This approach places high demands on the dataset and is often difficult to satisfy. Among commonly used datasets, we found that only the MIND dataset provides both unclicked items that appeared alongside the clicked ones and side information for those items. Additionally, some less common datasets, such as the ZhihuRec-Dataset, provide unclicked items but lack detailed side information, offering only token vectors. Furthermore, while classic CTR datasets include labels indicating whether a user clicked on an item, they are not suitable for our purpose, as their features are anonymous and do not contain textual information.
To address your concerns, we have relaxed the stringent requirements for the offline experiments on the dataset and included two additional datasets for validation. We consider items that are interacted with within a time window to be simultaneously recommended to a user. Items with a rating greater than or equal to 4 were considered clicked, while those with a rating below 4 were considered unclicked. We conducted experiments on the ML-1M and Amazon Book datasets, and the results were consistent with those obtained from the MIND dataset.
@@ F.0.12. DiscomfortFilter or DisinterestFilter?
F.0.13. The A/B test for the control group.
Our findings revealed that DiscomfortFilter’s design inherently accounts for contestability, ensuring transparency and user control over its decisions. For instance, users can review filtering log and understand the rationale behind each filtering record, with the ability to make adjustments if the filtering records do not meet their expectations. Consequently, when we presented users with a mock filtering interface—where no actual or only random filtering occurred—they quickly discerned they were part of the control group, leading to immediate expressions of dissatisfaction. Moreover, during the user study in current version, participants expressed satisfaction with the performance of the functionalities, not just their design (Section 5.2.4 & 5.2.5). In summary, we posit that the mere presence of a filtering interface is unlikely to enhance user satisfaction. Instead, satisfaction levels are significantly influenced by the efficacy of the actual filtering performance.

### v3

@@ Abstract.
discomforting recommendation filtering, large language model
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714850isbn: 979-8-4007-1274-6/25/04ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. The workflow of DiscomfortFilter.
Figure 2. Problem formulation.
Personalized algorithms, which analyze user preferences to deliver tailored content and thereby support human decision-making, are indispensable across web platforms (Xia et al., 2022; Liu et al., 2023b, d, 2022, a, c). While these algorithms are designed to enhance user experience, they can inadvertently expose users to discomforting recommendations (Park et al., 2024). For example, if a user searches for sensitive topics like health issues, the algorithm might suggest related health products, which could be perceived as a breach of privacy and cause unease (Ur et al., 2012). Similarly, when a user is experiencing emotional distress, such as after a breakup, algorithms lacking contextual awareness might recommend content that evokes painful memories, potentially worsening the user’s emotional state (Pinter et al., 2019). Such recommendations may not only fail to engage users but also lead to negative emotional consequences, such as anxiety, unease, or distress (Park et al., 2024). The perception of discomfort is highly subjective, meaning that content one user finds enjoyable may be discomforting to another (Park et al., 2024; Smith et al., 2022; Pierce et al., 2018; Singh, 2019). This subjectivity underscores the urgent need for a more nuanced approach to discomforting content identification that aligns more closely with individual user experiences (Harris and Cook, 2011; Reichmann et al., 2023; Shahi and Kana Tsoplefack, 2022).
@@ 2. Related Work
Personalized algorithms primarily derive user preferences from behavioral data, leading to incomplete user modeling and discomforting recommendations (Ricci et al., 2021). In Appendix E, we provide a detailed review of studies illustrating this phenomenon across various scenarios, including privacy invasion (Ur et al., 2012), lack of contextual understanding (Pinter et al., 2019), popularity bias (Ferraro, 2019), and information bubbles (Rowland, 2011), along with their negative outcomes (Smith et al., 2022). Although previous research has identified scenarios and causes of discomfort, few studies have focused on designing systems to identify and filter these recommendations, which our work aims to address.
In Appendix E, we provide a detailed review of two potential solutions for filtering discomforting recommendations: interactive recommendation systems (Lyons et al., 2021; He et al., 2016; Ge et al., 2022) and content moderation systems (Cresci et al., 2022; Gurkan et al., 2024; Jhaver et al., 2023). Both allow users to modify recommendations to mitigate discomfort. However, interactive systems often struggle to scale in opaque environments due to their reliance on specific algorithm designs, while moderation systems focus on objectively harmful content, which may be less effective for addressing subjective discomfort. Thus, the subjectivity of discomfort perception and the opacity of algorithms present notable challenges.
In Appendix E, we present a detailed review of studies on LLMs as personal assistants (Li et al., 2024), including commercial applications (Mehdi, 2023a) and frameworks for human-centered recommender systems (Shu et al., 2024). The impressive capabilities of LLMs, especially in handling personal data and services, underscore their potential for effectively filtering discomforting recommendations. To our knowledge, this is the first exploration of using LLMs for this specific purpose.
@@ 4.2.1. Content Filter Module
A user can manage filtering rules directly through this module. These rules are described in natural language, specifying the discomforting recommendations the user wishes to avoid. Each filtering rule has an associated activation option. During the filtering phase, the module reviews each recommendation against all active filtering rules to determine if the content matches any discomforting criteria. Only recommendations that do not match any filtering rules will be shown to the user; otherwise, they will be filtered out (G4). The identification process uses a chain-of-thought (CoT) method, with the prompt detailed in Appendix C. Filtering records will be saved and forwarded to the Filtering Needs Discovery Module.
@@ 4.2.2. Preference Profile Construction Module
The preference profile is stored and sent to the Filtering Needs Discovery Module to improve the personalization of the conversational agent and provide users with clear explanations. The prompt used for this multi-agent pipeline can be found in Appendix C.
@@ 5.1.1. Task
# Messages # Add # Update
Content Filter Module - 2.9 2.1
Strategy 1 13.0 4.4 2.7
Strategy 2 22.3 3.2 7.3
@@ 5.1.4. Results
Table 2. Confusion matrix.
Predicted Positive Predicted Negative
Actual Positive 173 32
Actual Negative 67 208
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (NSFC) under the Grant No. 61932007, 62372113, and 62172106. Tun Lu is also a faculty of Shanghai Key Laboratory of Data Science, Fudan Institute on Aging, MOE Laboratory for National Development and Intelligent Governance, and Shanghai Institute of Intelligent Electronics & Systems, Fudan University.
@@ E.2.2. Content Moderation Systems
Content moderation is used to regulate harmful content, such as hate speech and abuse (Alghowinem, 2019; Liu et al., 2023e; Franco et al., 2023; Jhaver et al., 2019; Ye et al., 2023; Park et al., 2021). The process has evolved from manual review (Gillespie, 2018; Seering et al., 2019) to automated methods using NLP technologies (Chandrasekharan et al., 2019; Horta Ribeiro et al., 2023; Long et al., 2017), which has led to reduced transparency in moderation and less feedback for users (Wright et al., 2021; Jhaver et al., 2019). To address these challenges, RECAST (Wright et al., 2021) provides mechanisms for explanation, revision, and user feedback in the moderation process. Recent research on using LLMs for content moderation (Ma et al., 2023; Markov et al., 2023; Ye et al., 2023; Mullick et al., 2023) highlights their potential in tasks like toxicity detection (Kumar et al., 2023) and rule violation identification (Kolla et al., 2024). At the same time, there is increasing support for decentralizing content moderation, shifting control from centralized platforms to individual users, thus better enabling them to manage content they wish to avoid on social media (Jhaver et al., 2022). This shift recognizes that a one-size-fits-all approach to content regulation fails to meet the diverse needs of the user base (Jiang et al., 2023; Jhaver et al., 2023).
@@ E.3. LLMs as Personal Assistants
Several commercial products, such as Microsoft’s Copilot, New Bing, Google’s Bard, and Gemini, as well as offerings from smartphone manufacturers, have integrated LLMs to boost productivity and enhance web and on-device experiences (Mehdi, 2023a; Achiam et al., 2023; Mehdi, 2023b). In the area of human-centered recommender systems, the RAH framework (Shu et al., 2024), an LLM-based multi-agent system, analyzes user preferences by observing ratings and reviews and incorporates a reflection mechanism to improve the accuracy.
@@ F.0.1. Generalizability to non-text-based items
Expanding to non-text scenarios is a natural progression, requiring the use of multimodal models to understand such content. However, the efficiency of processing, especially for videos, remains a concern. Preprocessing non-text content offline may offer a viable solution.
@@ F.0.2. Differences between discomforting and disliked items
F.0.3. Discussion of critiquing-based recommender systems
@@
F.0.4. Computational overhead and scalability
@@
F.0.5. Discussion on the echo chamber problem
@@
F.0.6. Comparison with recommendation unlearning
Recommendation unlearning emphasizes the model’s capability, with a primary focus on the model’s completeness, efficiency, and performance after forgetting. These methods typically predefine which interactions should be forgotten, assess the completeness of forgetting through Membership Inference Attacks (MIA), measure the efficiency of forgetting through runtime analysis, and evaluate post-forgetting performance through recommendation accuracy. In contrast, DiscomfortFilter filters interactions from the perspective of user needs—due to its subjectivity, it is difficult to clearly define which interactions should be filtered in advance. Therefore, it is hard to make a quantitative comparison with recommendation unlearning, as current unlearning methods cannot handle subjective user inputs. We believe that these two types of work focus on model design and human-centered design, respectively, and are complementary to each other.
F.0.7. The effect of ephemeral filtering criteria
@@
F.0.8. Only the MIND dataset was used in the offline experiment
DiscomfortFilter constructs positive and negative sample pairs by observing the content recommended by the algorithm and the user’s click behavior on that content. Specifically, our negative sampling is not performed randomly from the entire item set; instead, it is selected from items that appeared alongside the clicked content but were not clicked. This approach places high demands on the dataset and is often difficult to satisfy. Among commonly used datasets, we found that only the MIND dataset provides both unclicked items that appeared alongside the clicked ones and side information for those items. Additionally, some less common datasets, such as the ZhihuRec-Dataset, provide unclicked items but lack detailed side information, offering only token vectors. Furthermore, while classic CTR datasets include labels indicating whether a user clicked on an item, they are not suitable for our purpose, as their features are anonymous and do not contain textual information.
@@ F.0.12. DiscomfortFilter or DisinterestFilter?
F.0.13. The A/B test for the control group
Our findings revealed that DiscomfortFilter’s design inherently accounts for contestability, ensuring transparency and user control over its decisions. For instance, users can review filtering log and understand the rationale behind each filtering record, with the ability to make adjustments if the filtering records do not meet their expectations. Consequently, when we presented users with a mock filtering interface—where no actual or only random filtering occurred—they quickly discerned they were part of the control group, leading to immediate expressions of dissatisfaction. Moreover, during the user study in current version, participants expressed satisfaction with the performance of the functionalities, not just their design. In summary, we posit that the mere presence of a filtering interface is unlikely to enhance user satisfaction. Instead, satisfaction levels are significantly influenced by the efficacy of the actual filtering performance.